# Architectural Patterns for Designing Quantum Artificial Intelligence Systems

### v1

@@ 6 Discussion
The number of publications corresponding to each of the architectural patterns described above is shown in Fig. 11). The most widely used pattern is the quantum monolith. This architecture is the simplest and is well-suited for demonstrating proof-of-concepts and conducting experiments in the field of QML. The next most used patterns are quantum head and intermediate quantum layer. These architectural solutions are closely tied to real-world AI applications for NISQ-era quantum computers. They are well-suited for high-dimensional data and offer promising advantages in the near future.
Uploading large amounts of data directly into quantum computers poses challenges for NISQ-era systems due to limitations in the number of available qubits and their coherence time. As a result, the quantum feature extraction pattern is less commonly used. This pattern is mostly utilized in quantum support vector machines, where the quantum component operates directly on data to compute the feature map. The limitations related to a small number of qubits are somewhat alleviated by quantum convolutional filters (Quanvolution), where the filter scans over the input data layer, processing only a small amount of data at a time. These systems are expected to significantly reduce the number of learnable parameters without loss of expressibility. However, this potential has not yet been fully realized on NISQ computers since it requires deep integration and communication between classical and quantum components.
@@
Further advancements in quantum architecture search include a cloud-based automated circuit architecture search framework proposed in Berganza Gómez et al. (2022). This framework enables parallelized hyperparameter exploration and model training, leveraging the power of cloud computing. Recognizing the critical role of benchmarking in advancing QAS, Lu et al. developed QAS-bench, a benchmark framework comprising two datasets with 900 quantum circuits and 400 unitary matrices, along with evaluation protocols to assess both existing and future QAS approaches Lu et al. (2023b).
@@
Distributed quantum computing Distributed quantum computing presents an architecturally scalable solution for building large-scale quantum computing systems, addressing the limitations of NISQ systems. For example, a study explored distributing partitioned features across multiple smaller, spatially separated quantum circuits (nodes), resulting in a slight decrease in performance compared to a monolithic quantum neural network Wu et al. (2023); Kawase (2024). While dataset distribution shares similarities with classical approaches, the distribution of quantum models is distinct due to the need for distributed quantum networks to establish entanglement across different nodes Pira and Ferrie (2023). The inter-node two-qubit gates can be managed using three approaches: terminating two-qubit gates with measurements, approximating them with local gates, or establishing a coherent quantum communication channel between non-local qubits Pira and Ferrie (2023).
Physically, the latter approach involves long-range communication with fiber optics to establish entangled Einstein-Podolsky-Rosen photon pairs Häner et al. (2021); Wu et al. (2023). Communication between nodes cannot be broadband; thus, optimal partitioning of quantum circuits to minimise inter-node two-qubit gates is essential. This optimisation can be automated at the compiler level, as implemented in the QuComm framework Wu et al. (2023), or manually using the Quantum Message Passing Interface, analogous to MPI functions in multi-processing programming Häner et al. (2021). Additionally, the concept of federated learning, which has gained significant attention in classical deep learning, is considered beneficial for quantum machine learning McMahan et al. (2017); Li et al. (2021c); Larasati et al. (2022); Xia and Li (2021); Chehimi and Saad (2022).
@@
7 Related works
Quantum software engineering, an emerging subfield of software engineering, aims to apply the principles and practices of software design and development to the creation of quantum algorithms that can be executed or simulated on quantum systems. At a high level, quantum software is developed using extensions of scripting programming languages such as Q#, Qiskit, Cirq, Quipper, and Scaffold Heim et al. (2020). The developed program is then transpiled into an intermediate low-level representation, typically in the language OpenQASM, which serves as a gate-based intermediate representation for quantum programs Bandic et al. (2022). Transpilation involves rewriting a given input circuit to match the topology of a specific quantum device and/or optimizing the circuit for execution on current noisy quantum systems. This intermediate representation is subsequently compiled into a specific instruction set architecture tailored for dedicated hardware. The quantum software development lifecycle shares many similarities with classical software development Zhao (2020), although it involves certain unique aspects as described in Weder et al. (2020a, 2022); Dwivedi et al. (2024).
Significant effort has been invested in identifying architectural and design patterns in quantum software Leymann (2019); Weigold et al. (2021b, a); Georg et al. (2023); Khan et al. (2022), including the proposal of a specialized pattern language Leymann (2019). The aforementioned works address quantum software in general. In this paper, we focus on architectural patterns that are specific to quantum AI and quantum machine learning.
@@ 9 Conclusions
The unique software architectural solutions concerning quantum AI primarily revolve around determining which tasks to delegate to a quantum computer and how quantum and classical software components communicate with each other. The primary solution for the latter is using a service-oriented architecture. During the systematic mapping study, we identified ten architectural patterns in quantum AI systems: seven related to the quantum-classic split and three related to quantum middleware. Each architectural pattern involves trade-offs in achieving the system quality attributes, such as efficiency, scalability, trainability, simplicity, portability, and deployability. Balancing these trade-offs is a common challenge in software engineering. These patterns constitute the core of the pattern catalog developed in this work, which aims to guide software architects in effectively utilising quantum components in their projects.
In quantum AI systems, the use of quantum components can be categorised into two main areas: quantum as part of trainable inference models and quantum as an accelerator for specific functions. Quantum AI systems offer several benefits, such as accelerating the inference and training stages, enhancing robustness against noise in data and adversarial attacks, and achieving better accuracy. However, most of the collected evidence confirming these advantages is obtained from simulations on classical computers. There is no evidence that NISQ-era quantum computers can provide these benefits for practical applications. The identified trends in quantum AI are related to scaling up quantum components, automatic quantum architecture search, hardware-software co-design, utilising distributed quantum computing and federated learning, and developing quantum DevOps.

### v2

@@ 6 Discussion
The number of publications corresponding to each of the architectural patterns described above is shown in Fig. 11. The most widely used pattern is the quantum monolith. This architecture is the simplest and is well-suited for demonstrating proof-of-concepts and conducting experiments in the field of QML. The next most used patterns are quantum head and intermediate quantum layer. These architectural solutions are closely tied to real-world AI applications for NISQ-era quantum computers. They are well-suited for high-dimensional data and offer promising advantages in the near future.
Uploading large amounts of data directly into quantum computers poses challenges for NISQ-era systems due to limitations in the number of available qubits and their coherence time. As a result, the quantum feature extraction pattern is less commonly used. This pattern is mostly utilised in quantum support vector machines, where the quantum component operates directly on data to compute the feature map. The limitations related to a small number of qubits are somewhat alleviated by quantum convolutional filters (Quanvolution), where the filter scans over the input data layer, processing only a small amount of data at a time. These systems are expected to significantly reduce the number of learnable parameters without loss of expressibility. However, this potential has not yet been fully realised on NISQ computers since it requires deep integration and communication between classical and quantum components.
@@
Further advancements in quantum architecture search include a cloud-based automated circuit architecture search framework proposed in Berganza Gómez et al. (2022). This framework enables parallelised hyperparameter exploration and model training, leveraging the power of cloud computing. Recognizing the critical role of benchmarking in advancing QAS, Lu et al. developed QAS-bench, a benchmark framework comprising two datasets with 900 quantum circuits and 400 unitary matrices, along with evaluation protocols to assess both existing and future QAS approaches Lu et al. (2023b).
@@
Distributed quantum computing Distributed quantum computing presents an architecturally scalable solution for building large-scale quantum computing systems, addressing the limitations of NISQ systems. Several experiments have been conducted in this direction. Several authors Wu et al. (2023); Kawase (2024) explored the distribution of partitioned features across multiple spatially separated quantum circuits (nodes) and reported a slight decrease in performance compared to a monolithic quantum neural network. While dataset distribution shares similarities with classical approaches, the distribution of quantum models is distinct due to the need for distributed quantum networks to establish entanglement across different nodes Pira and Ferrie (2023). The inter-node two-qubit gates can be managed using three approaches: terminating two-qubit gates with measurements, approximating them with local gates, or establishing a coherent quantum communication channel between non-local qubits Pira and Ferrie (2023).
Physically, the latter approach involves long-range communication with fibre optics to establish entangled Einstein-Podolsky-Rosen photon pairs Häner et al. (2021); Wu et al. (2023). Communication between nodes cannot be broadband; thus, optimal partitioning of quantum circuits to minimise inter-node two-qubit gates is essential. This optimisation can be automated at the compiler level, as implemented in the QuComm framework Wu et al. (2023), or manually using the Quantum Message Passing Interface, analogous to MPI functions in multi-processing programming Häner et al. (2021). Additionally, the concept of federated learning, which has gained significant attention in classical deep learning, is considered beneficial for quantum machine learning McMahan et al. (2017); Li et al. (2021c); Larasati et al. (2022); Xia and Li (2021); Chehimi and Saad (2022).
@@ 7 Threats to validity
Another challenge discussed in Ronnow et al. (2014) concerns the experimental evaluation of quantum speedup as reported in the literature. The authors argue that, to properly distinguish parallel speedup from quantum speedup, it is crucial to scale hardware resources identically for both quantum and classical devices during comparisons. Moreover, it is important to avoid conflating inefficiencies at small problem sizes with signs of quantum speedup.
@@ 8 Conclusions
The unique software architectural solutions concerning quantum AI primarily revolve around determining which tasks to delegate to a quantum computer and how quantum and classical software components communicate with each other. During the systematic mapping study, we identified ten architectural patterns in quantum AI systems: seven related to the quantum-classic split and three related to quantum middleware. These patterns constitute the core of the pattern catalog developed in this work, which aims to guide software architects in effectively utilising quantum components in their projects. Each architectural pattern involves trade-offs in achieving the system quality attributes, such as efficiency, scalability, trainability, simplicity, portability, and deployability. The most commonly used patterns for small-scale proof-of-concept applications are the quantum monolith and quantum feature engineering. When attempting to scale the system to handle larger volumes of information within the constraints of NISQ-era hardware, the corresponding challenges are typically addressed using architectural patterns such as the quantum head or quantum intermediate layer. In these architectures, part of the inference task, as well as the reduction of data dimensionality, is delegated to classical components. The primary approach for organising communication between classical and quantum components is through the use of a service-oriented architecture. In quantum AI systems, quantum components can be integrated into trainable inference models or function as accelerators for specific tasks.
Quantum AI systems offer several benefits, such as accelerating the inference and training stages, enhancing robustness against noise in data and adversarial attacks, and achieving better accuracy. However, most of the collected evidence confirming these advantages is obtained from simulations on classical computers. There is no yet evidence that NISQ-era quantum computers can provide these benefits for practical applications. The identified trends in quantum AI are related to scaling up quantum components, automatic quantum architecture search, hardware-software co-design, utilising distributed quantum computing and federated learning, and developing quantum DevOps.